OpenStudy (anonymous):

cos2x maclarins series

OpenStudy (anonymous):

well you will have f(x)= cos2x then you need have f ' and f ''

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

ok sorry im reading my old calc book and relearning this

OpenStudy (anonymous):

the maclaurin series is generated by f. \[\sum_{k=0}^{\infty} (f ^{k}(0)/ k!)x ^{k} = f(0) + f'(0)x + (f''(0)/2!)(x ^{2})+...+...(f ^{n}(0)/n!) x ^{n}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

did you already have that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you want the answer and steps to get there

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

differential cos2x to get there

OpenStudy (anonymous):

f '= -2sin(2x) f ''=-4cos(2x)

OpenStudy (anonymous):

f3 as the next one

OpenStudy (anonymous):

f '''= 8sin(2x) f ''''=16cos(2x)

OpenStudy (anonymous):

sin(0)=0 cos(0)= 1

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

bare with me here im going to type all i have here

OpenStudy (anonymous):

-sin(0)= 0 -cos(o)=-1

OpenStudy (anonymous):

do cos2x differential it like 5 times

OpenStudy (anonymous):

cos(2(0)) -2sin(2(0))x -(4cos(2(0))/2)(x^2)+(8sin(2(0))/3!)(x^3)+(16cos(2(0))/4!)(x^4) -(32sin(2(0))/5!)(x^5) and i will simplify here in a sec

OpenStudy (anonymous):

1-0-2(x^2)+0+16(x^4)+0

OpenStudy (anonymous):

you follow me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this the answer you were looking for?